OpenStudy (anonymous):

Simplify the given expression. (3 + 4i) + (5 − 2i)

OpenStudy (anonymous):

first make the equation out of bracket 3+4i+5-2i then collect the like terms =2i+8 and if you have to solve for i then 2i=-8 i=-8/2 i=-4